OpenStudy (anonymous):

Write an equation for a circle with center at (-2,7) and diameter of 14

OpenStudy (aakashsudhakar):

(h, k) ---> (-2, 7) D = 14 D = 2r 2r = 14 r = 7 (x - h)^2 + (y - k)^2 = r^2 Plug in all values given. (x + 2)^2 + (y - 7)^2 = 49

OpenStudy (anonymous):

(X+2)^2 + (Y-7)^2 = 7^2 As genral equation of circle is (x-h)^2 + (Y-k)^2=r^2 where h cuts on x axis k cuts on y axis and r is radius= (diameter/2)

OpenStudy (anonymous):

wait so whats the answer 2?

OpenStudy (anonymous):

(X+2)^2 + (Y-7)^2 = 7^2

OpenStudy (anonymous):

oo alright sorry thanks

OpenStudy (anonymous):

or (X+2)^2 + (Y-7)^2 = 49

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

sorry my thing online said both were wrong ;(

OpenStudy (anonymous):

not possible

OpenStudy (anonymous):

\[\left( x+2 \right)^{2}+(y-7)^{2}=49\]

OpenStudy (aakashsudhakar):

^^ Like I said too, what @reshabhraj just said should be the correct answer.

OpenStudy (anonymous):

now it said its right...sorry guys my online class is dumb sometimes now it accepted it though thank yall